What is the probability that it is either blue or orange


1
Expert's answer
2021-05-18T07:04:26-0400

Actually compl question is :-


An urn contains 3 red, 5 blue, 4 green and 6 orange marbles. A marble is drawn from the urn at

random. What is the probability that it is:

(a) a green marble?

(b) not a red marble?

(c) either blue or orange marble?

(d) an orange marble?


Solution:-


Total marble = 18


(i) probability of a green marble

=green marble / total marble

=4/18

=2/9


(ii)probability of not a red marble

=1-(red marble / total marble)

=1-(3/18)

=5/6


(iii)probability of either blue or orange marble

=blue+orange marble / total marble

=11/18


(iv)probability of a orange marble

=orange marble / total marble

=6/18

=1/3